# Mpemba effect on non-equilibrium active Markov chains

### v1

@@ Abstract
We study the Mpemba effect on a non-equilibrium Markov chain that mimics the run and tumble motion of an active particle in a discrete energy landscape. The broken detailed balance, rendered by the activity, gives rise to a unique anomalous relaxation in the system which is distinctly different than the typical equilibrium systems. We observe that the activity can both suppress or induce the Mpemba effect. Furthermore, we report an oscillatory Mpemba effect where the relaxation trajectories, emanating from the hot and cold initial conditions, cross each other multiple times and this occurs due to the emergence of complex eigenvalues in the relaxation spectrum due to the activity. Our work reveals a possible pathway for studying the Mpemba effect in active living systems where broken detailed balance is crucial to achieve many biological functions.
@@
The Mpemba effect refers to the anomalous relaxation phenomena where an initially hotter system equilibrates faster than an initially cooler system, when both systems are quenched to the same low temperature [1]. Although the phenomena was first observed in the case of water namely, hot water freezing faster [1, 2, 3, 4, 5, 6, 7, 8, 9] than an equal volume of warm water exposed to the same condition, it is now a well-observed phenomena that has been probed through both theoretical and experimental means across various systems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 24, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. A particular frontier that remains notably unexplored is the effect of activity in the anomalous relaxation dynamics (besides the work by Schwarzendahl and Löwen on active colloids [29]). Activity is a fingerprint of non-equilibrium systems which unlike their equilibrium counterpart violate detailed balance and this results in many interesting features such as non-Boltzmannian steady states, clustering, spontaneous aggregation, motility induced phase transition in active self-propelled systems, sense and adaption for high fidelity DNA transcription, structural organization in cells, folding of macromolecules [53]. In this paper, we aim to understand how the broken detailed balance, specifically induced due to the activity, can manifest in the anomalous relaxation dynamics such as the Mpemba effect.
Figure 1: Illustration of an active Markov chain model capturing the dynamics of a run and tumble motion of an active walker in a discrete 3-state energy landscape. Panel (a): The hopping rates $R_{ij}$ that satisfy detailed balance and in effect, characterize the rates of diffusion of an active walker among the energy states. Panel (b) depicts the corresponding energy landscape where $B_{ij}$’s are the energy barriers between states with energies $E_{i}$ and $E_{j}$. Panel (c): The activity of the walker is characterised by the rates $\gamma$ and $\epsilon$, where $\gamma$ is the flipping rate between the states $\uparrow$ and $\downarrow$ – equivalent to the persistence between the two states, and $\epsilon$ is the state-dependent rate of unidirectional hopping that resembles the velocity of the active walker. In essence, $\gamma\neq 0$ is the non-equilibrium active component that drives the system away from equilibrium.
@@
In particular, we show via a rigorous analysis that an active system can display a rich behaviour in the relaxation dynamics such as the emergence of oscillations in the transient & the activity -induced or -suppressed Mpemba effect. The results are intriguing as the presence of such behaviours in the eigenspectrum or in the relaxation dynamics has no counterpart in regular equilibrium setups. Although a similar oscillatory behaviour in the relaxation has been reported recently in a two-level quantum system where such behaviour is attributed to the role played by the presence of exceptional points and complex eigenvalues [54], the presence of such behaviours in the classical models is rather unique which we explore herein. Finally, the presence of non-equilibrium steady states and the oscillatory relaxations raises plausible questions on the adaptable measures (pertaining only to the truncated eigenspectrum) that are often used to quantify the Mpemba effect in equilibrium systems [27, 22].
@@
Model and framework: Consider a setup of a discrete three state active Markov chain (Fig. 1). The intrinsic hopping rate of a random walker from the state $j$ to $i$ is denoted by $R_{ij}$ and vice-versa. These rates satisfy detailed balance and are determined at a fixed bath temperature $T_{b}$ by the following relation
@@
where $E_{i}$ are the energy levels of the three states and $B_{ij}$ are the barrier heights between any two states $i$ and $j$. These rates usually characterize the rates of diffusion of the walker among the different energy states. To introduce the activity, we assume that there is an internal spin degree of freedom (denoted by $\uparrow$ and $\downarrow$) associated with each state. The spin can flip with a rate $\gamma$ which characterizes the persistence of the walker to be in that particular internal state. Depending on its current internal state ($\uparrow$ or $\downarrow$), the walker hops either to the clockwise or the anti-clockwise site at a rate $\epsilon$, thus characterizing the velocity of the active walker [55]. Hence, this model mimics a prototypical run and tumble motion of an active walker [55] on a discrete energy landscape.
@@
and similarly, for the sites $i=2,3$. Here, $\bar{\sigma}$ denotes the conjugate of $\sigma$. To solve for the occupation probability, it is useful to define the following quantities
@@
where the vector $\boldsymbol{P}=(P_{1},P_{2},P_{3},P^{*}_{1},P^{*}_{2},P^{*}_{3})^{\mathbb{T}}$ denotes the instantaneous probability distributions of the system at time $t$ and $\boldsymbol{W}(T_{b})$ is the transition matrix that is determined at the bath temperature $T_{b}$. Here, the initial condition is chosen to be a steady state probability distribution $\boldsymbol{P}(T,t=0)$ characterized by a different bath temperature $T\neq T_{b}$ to which the system is initially prepared before being quenched to the final bath temperature $T_{b}$. If the transitions between microscopic states are pairwise-balanced (like in Eq. (1)), the system reaches a thermodynamic equilibrium precluding net flux among the states. On the other hand, if the transitions are not balanced, the system can exhibit flux loops, and moreover attains a non-equilibrium steady state.
Figure 2: Eigenvalues $\{\lambda\}$-s of the transition matrix $\boldsymbol{W}$ of an active Markov chain (with broken detailed balance): Presence of both real and complex eigenvalues is observed as we change the activity parameters. The transition from real to mixed eigenvalues as $\gamma$ is varied for fixed $\epsilon=1.0$ [(in panel (a)] and $\epsilon=9.0$ [(in panel (b)].
@@
The eigenvalues follow the order $\lambda_{1}=0>\text{Re}\lambda_{2}\geq\text{Re}\lambda_{3}\ldots,$ where $\lambda_{1}=0$ is the largest eigenvalue that corresponds to the eigenvector $\boldsymbol{\pi}(T_{b})$. All other eigenvalues have negative real parts. The coefficient $a_{i}(T,T_{b})$ in Eq. (5) is the overlap coefficient of the $i^{th}$ eigenvector with the initial distribution $\boldsymbol{P}(T,t=0)$.
@@
If the transition rates satisfy detailed balance, it is guaranteed that all the eigenvalues are real [56]. On the other hand, for non-equilibrium dynamics with broken detailed balance, the eigenspecturm can be mixed with the presence of both real and complex eigenvalues. Here too, we observe that the eigenvalues turn out to be complex as one varies the activity parameter $\gamma$. Figure 2 illustrates such a transition in the eigenspectrum as a result of the change in the active parameters $\epsilon$ and $\gamma$.
@@
Figure 3: Activity induced and suppressed Mpemba effect. Panel (a): The Mpemba effect is absent in the passive case ($\epsilon=\gamma=0$) while the presence of activity ($\epsilon=2$ and $\gamma=4$) induces the Mpemba effect (shown in inset). The model parameters are: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.4$, $B_{13}=B_{31}=0.5$, $B_{23}=B_{32}=0.4$, $T_{h}=4.0$, $T_{c}=1.0$ and $T_{b}=0.1$. Panel (b) shows activity suppressed Mpemba effect (shown in inset) while it is present in the passive case. Here, the active rates are $\epsilon=1$ and $\gamma=2$. The model parameters in this case are: $E_{1}=0.2$, $E_{2}=0.3$, $E_{3}=0.5$, $B_{12}=B_{21}=0.9$, $B_{13}=B_{31}=0.6$, $B_{23}=B_{32}=0.9$, $T_{h}=3.0$, $T_{c}=1.0$ and $T_{b}=0.1$. Although the measure $L_{1}$-norm has been used as a reference, the results are found to be independent of the distance measures [SM].
@@
The Mpemba effect: To interpret the Mpemba effect, let us consider two systems: first one $A$, initially in a steady state corresponding to a hot temperature $T_{h}$ and second one $B$, initially in a different steady state corresponding to cold temperature $T_{c}$ where $T_{h}>T_{c}$. These initial steady states are denoted by $\boldsymbol{\pi}(T_{h})$ and $\boldsymbol{\pi}(T_{c})$ respectively. Now imagine that both $A$ and $B$ are quenched at once to a common bath temperature, $T_{b}$, where $T_{h}>T_{c}>T_{b}$. Eventually, both of them will relax to the common distribution $\boldsymbol{\pi}(T_{b})$ given long enough time. The Mpemba effect is said to exist if $A$ reaches the final steady state faster than $B$ during the transient/relaxation process. This relaxation can be quantified in terms of the distance function $D[\boldsymbol{\mathcal{P}}(t),\boldsymbol{\Pi}(T_{b})]$ which measures the instantaneous distance of the state-wise occupation probabilities $\boldsymbol{\mathcal{P}}(t)=(P_{1},P_{2},P_{3})^{\mathbb{T}}$ from their respective final steady state distributions $\boldsymbol{\Pi}(T_{b})=(\pi_{1},\pi_{2},\pi_{3})^{\mathbb{T}}$. There are many well-adapted measures that exist in the literature namely the norm $L_{1}(t)=\sum_{i=1}^{3}|P_{i}(t)-\pi_{i}(T_{b})|$ and Kullback-Leibler (KL) divergence $KL(t)=\sum_{i=1}^{3}P_{i}(t)\ln(P_{i}(t)/\pi_{i})$. Thus, if one has $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]>D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ initially for $T_{h}>T_{c}$ followed by $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]<D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ at a later time, we infer that the Mpemba effect exists.
Activity harnessed Mpemba effect: To highlight the effect of activity, we first set $\epsilon=\gamma=0$ and consider the passive discrete three-state Markov chain [27]. The identification of the Mpemba effect in this case can be made from the behaviour of the coefficient $|a_{2}(T,T_{b})|$ [as in Eq. (5)] [27]. In particular, we choose a set of parameters of the system (namely $R_{ij}$-s) where the Mpemba effect is absent. With the present parameters as the activity is introduced, we find an emergence of the Mpemba effect [see Fig. 3]. We also note a converse effect where the introduction of activity suppresses an already existing Mpemba effect [see Fig. 3 for more details]. Notably, this behavior remains quite robust under other distance measures [see SM].
Figure 4: Illustration of the oscillatory Mpemba effect across different distance measures: (a) the $L_{1}$-norm, and (b) the Kullback-Leibler (KL) divergence for a system with broken detailed balance. The complex eigenspecturm can lead to oscillatory relaxations and distance measures. Here, the model parameters are: $E_{1}=0.0$, $E_{2}=0.1$, $E_{3}=0.3$, $B_{12}=B_{21}=0.4$, $B_{13}=B_{31}=0.5$, $B_{23}=B_{32}=0.4$, $\gamma=8$, $\epsilon=9$ whereas the temperatures of the hot, cold and the bath temperatures are chosen to be: $T_{h}=5.0$, $T_{c}=0.5$ and $T_{b}=0.1$, respectively.

### v2

@@ Abstract
We study the Mpemba effect on a non-equilibrium Markov chain that imitates the active motion of random walkers in a discrete energy landscape. The broken detailed balance, rendered by the activity, gives rise to a unique anomalous relaxation in the system which is distinctly different than the typical equilibrium systems. We observe that the activity can both suppress or induce the Mpemba effect. Furthermore, we report an oscillatory Mpemba effect where the relaxation trajectories, emanating from the hot and cold initial conditions, cross each other multiple times and this occurs due to the emergence of complex eigenvalues in the relaxation spectrum due to the activity. Our work reveals a possible pathway for studying the Mpemba effect in active living systems where broken detailed balance is crucial to achieve many biological functions.
I Introduction
@@
The Mpemba effect refers to the anomalous relaxation phenomena where an initially hotter system equilibrates faster than an initially cooler system, when both systems are quenched to the same low temperature [1]. Although the phenomena was first observed in the case of water namely, hot water freezing faster [1, 2] than an equal volume of warm water exposed to the same condition, it is now a well-observed phenomena that has been probed through both theoretical and experimental means across various systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. A particular frontier that remains notably unexplored is the effect of activity in the anomalous relaxation dynamics (besides the work by Schwarzendahl and Löwen on active colloids [17]). Activity is a hallmark feature of non-equilibrium systems which unlike their equilibrium counterpart violate detailed balance and this results in many interesting features such as non-Boltzmannian steady states, clustering, spontaneous aggregation, motility induced phase transition in active self-propelled systems, sense and adaption for high fidelity DNA transcription, structural organization in cells, folding of macromolecules [34]. In this paper, we aim to understand how the broken detailed balance, specifically induced due to the activity, can manifest in the anomalous relaxation dynamics such as the Mpemba effect.
@@
In particular, we show via a rigorous analysis that an active system can display a rich behaviour in the relaxation dynamics such as the emergence of oscillations in the transient and the activity -induced or -suppressed Mpemba effect. The results are intriguing as the presence of such behaviours in the eigenspectrum or in the relaxation dynamics has no counterpart in regular equilibrium setups. Although a similar oscillatory behaviour in the relaxation has been reported recently in a two-level quantum system where such behaviour is attributed to the presence of exceptional points [35], the presence of such behaviours in the classical models is rather unique which we explore herein. Finally, the emergence of non-equilibrium steady states and the oscillatory relaxations raises plausible questions on the adaptable measures (pertaining only to the truncated eigenspectrum) that are often used to quantify the Mpemba effect in equilibrium systems [12, 15].
Figure 1: Illustration of an active Markov chain model capturing the dynamics of a run and tumble motion of an active walker in a discrete 3-state energy landscape. Panel (a): The hopping rates $R_{ij}$ that satisfy detailed balance and in effect, characterize the rates of diffusion of an active walker among the energy states. Panel (b) depicts the corresponding energy landscape where $B_{ij}$’s are the energy barriers between states with energies $E_{i}$ and $E_{j}$. Panel (c): The activity of the walker is characterized by the rates $\gamma$ and $\epsilon$, where $\gamma$ is the flipping rate between the states $\uparrow$ and $\downarrow$ – equivalent to the persistence between the two states, and $\epsilon$ is the state-dependent rate of unidirectional hopping that resembles the velocity of the active walker. In essence, $\gamma\neq 0$ is the non-equilibrium active component that drives the system away from equilibrium.
The remainder of the paper is organized as follows. Section II contains the model and its framework. The general definition of the Mpemba effect and the criteria for its occurrence are discussed in Sec. III. Next, we describe the effect of activity on the Mpemba effect along with qualitative reasoning in Sec. IV. Section V describes the oscillatory Mpemba effect and the effect of activity on its phase diagram in terms of the model parameters, and a comparison between the oscillatory Mpemba effect and a damped harmonic oscillator is discussed in Sec. VI. Section VII illustrates the possibility of having multiple crossings of trajectories even in the presence of real eigenspectrum and Sec. VIII contains the summary of results and a discussion of their implications.
II Model and framework
@@
Consider a setup of a discrete three state active Markov chain (Fig. 1). The intrinsic hopping rate of a random walker from the state $j$ to $i$ is denoted by $R_{ij}$ and vice-versa. These rates satisfy detailed balance and are determined at a fixed bath temperature $T_{b}$ by the following relation
@@
where $E_{i}$ are the energy levels of the three states, $B_{ij}$ are the barrier heights between any two states $i$ and $j$ and $k_{B}$ is Boltzmann constant. These rates usually characterize the rates of diffusion of the walker among the different energy states. To introduce the activity, we assume that there is an internal spin degree of freedom (denoted by $\uparrow$ and $\downarrow$) associated with each state. The spin can flip with a rate $\gamma$ which characterizes the persistence of the walker to be in that particular internal state. Depending on its current internal state ($\uparrow$ or $\downarrow$), the walker is biased to hop either clockwise, i.e., in the direction of $1\rightarrow 2\rightarrow 3\rightarrow 1$ for $\uparrow$ state or anti-clockwise for $\downarrow$ state at a rate $\epsilon$, thus characterizing the velocity of the active walker [36]. Notably, our system stems from the active lattice gas models with Ising interaction that were proposed to describe flocking [37, 38]. The Markov chain kinetic model used here are also fundamental to prototypical unimolecular enzyme reaction cycles or cellular biochemical switches that can be out of chemical equilibrium due to chemical-potential difference and consequently, the rates will not be detailed balanced (see [39] for an extensive review on this topic). Note that the breakdown of detailed balance is explicit in our system as the diffusive rates are energy balanced while the rates for flipping and directional hopping are not. In our analysis, we will be using the dimensionless variables. The time is measured in units of $\epsilon^{-1}$, the energy levels and barrier heights are measured in units of $k_{B}T_{b}$ but we set $k_{B}=1$ for simplicity. Any initial temperature $T$ is measured in units of the final bath temperature $T_{b}$, and the activity rates $\epsilon$ and $\gamma$ are measured in units of the intrinsic rate $R_{12}(T_{b})$.
@@
and similarly, for the sites $i=2,3$ (see Appendix A for details). Here, $\bar{\sigma}$ denotes the conjugate of $\sigma$. To solve for the occupation probability, it is useful to define the following quantities
@@
where the vector $\boldsymbol{P}=(P_{1},P_{2},P_{3},P^{*}_{1},P^{*}_{2},P^{*}_{3})^{\mathbb{T}}$ denotes the instantaneous probability distributions of the system at time $t$ and $\boldsymbol{W}(T_{b})$ is the transition matrix that is determined at the bath temperature $T_{b}$ [the form for $\boldsymbol{W}(T_{b})$ is given in Eq. (24)]. Here, the initial condition is chosen to be a steady state probability distribution $\boldsymbol{P}(T,t=0)$ characterized by a different bath temperature $T\neq T_{b}$ to which the system is initially prepared before being quenched to the final bath temperature $T_{b}$. If the transitions between microscopic states are pairwise-balanced [like in Eq. (1)], the system reaches a thermodynamic equilibrium precluding net flux among the states. On the other hand, if the transitions are not balanced, the system can exhibit flux loops, and moreover attains a non-equilibrium steady state.
Figure 2: Eigenvalues $\{\lambda\}$-s of the transition matrix $\boldsymbol{W}$ of an active Markov chain (with broken detailed balance): Presence of both real and complex eigenvalues is observed as we change the activity parameters. The transition from real to mixed eigenvalues as $\gamma$ is varied for fixed $\epsilon/R_{12}(T_{b})=20.09$ [(in panel (a)]. Panel (b) explores the behavior in the number of real eigenvalues, from two (minimum) to six (maximum), as a function of $\epsilon$ and $\gamma$. The other relevant parameters are set as in Figs. 5 and 7.
@@
The eigenvalues follow the order $\lambda_{1}=0>\text{Re}\lambda_{2}\geq\text{Re}\lambda_{3}\ldots,$ where $\lambda_{1}=0$ is the largest eigenvalue that corresponds to the eigenvector $\boldsymbol{\pi}(T_{b})$. In Eq. (5), the coefficient $a_{i}(T,T_{b})=\boldsymbol{v}^{\prime}_{i}.\boldsymbol{X}^{-1}\boldsymbol{P}(T,t=0)/\boldsymbol{v}^{\prime}_{i}.\boldsymbol{v}^{\prime}_{i}$ is the overlap coefficient of the $i^{th}$ eigenvector $\boldsymbol{v}^{\prime}_{i}$ with the initial distribution $\boldsymbol{P}(T,t=0)$. Here, $\boldsymbol{X}$ is formed from the eigenvectors $\boldsymbol{v}$ of $\boldsymbol{W}$ and the transformed eigenvectors $\boldsymbol{v}^{\prime}$ are obtained using $\boldsymbol{v}^{\prime}=\boldsymbol{X}^{-1}\boldsymbol{v}$ that are mutually orthogonal (see Appendix C for details). Note that since the transition matrix, $\boldsymbol{W}$ is in general non-Hermitian due to the broken detailed balance, the expansion in Eq. (5) is possible only in the case when its eigenvectors $\boldsymbol{v}$’s are mutually independent of each other. The criterion to check that its eigenvectors are independent is that the determinant of the matrix $\boldsymbol{X}$ (see Appendix C) formed by taking its columns as the eigenvectors $\boldsymbol{v}$ of $\boldsymbol{W}$ is non-zero. If there is a case where the determinant is zero, then the eigenvectors $\boldsymbol{v}$ do not form a basis. In that scenario, matrix $\boldsymbol{W}$ cannot be diagonalized and hence orthogonal conjugate eigenvectors $\boldsymbol{v}^{\prime}$ also do not exist.
@@
If the transition rates satisfy detailed balance, it is guaranteed that all the eigenvalues are real [40]. On the other hand, for non-equilibrium dynamics with broken detailed balance, the eigenspecturm can be mixed with the presence of both real and complex eigenvalues (see Appendix B). Here too, we observe that the eigenvalues turn out to be complex as one varies the activity parameter $\gamma$. Figure 2 illustrates such a transition in the eigenspectrum as a result of the change in the active parameters $\epsilon$ and $\gamma$. Figure 2(b) further describes the pattern in the eigenspectrum as $\epsilon$ and $\gamma$ are varied. The eigenspectrum is purely real for the combination of very small $\epsilon$ while fixing $\gamma$ either small or large. On the contrary, the minimum number (two) of real eigenvalues occurs at larger $\epsilon$, for a given $\gamma$.
@@
Figure 3: Activity induced and suppressed Mpemba effect. Panel (a): The Mpemba effect is absent in the passive case ($\epsilon/R_{12}(T_{b})=\gamma/R_{12}(T_{b})=0$) while the presence of activity ($\epsilon/R_{12}(T_{b})=40.17$ and $\gamma/R_{12}(T_{b})=80.34$) induces the Mpemba effect (shown in inset). The model parameters are: $E_{1}=0$, $E_{2}=T_{b}$, $E_{3}=3T_{b}$, $B_{12}=B_{21}=4T_{b}$, $B_{13}=B_{31}=5T_{b}$, $B_{23}=B_{32}=4T_{b}$, $T_{b}=0.1$, $T_{h}/T_{b}=40$ and $T_{c}/T_{b}=10$. Panel (b) shows activity suppressed Mpemba effect (shown in inset) while it is present in the passive case. Here, the active rates are $\epsilon/R_{12}(T_{b})=403.43$ and $\gamma/R_{12}(T_{b})=806.86$. The model parameters in this case are: $E_{1}=2T_{b}$, $E_{2}=3T_{b}$, $E_{3}=5T_{b}$, $B_{12}=B_{21}=9T_{b}$, $B_{13}=B_{31}=6T_{b}$, $B_{23}=B_{32}=9T_{b}$, $T_{b}=0.1$, $T_{h}/T_{b}=30$ and $T_{c}/T_{b}=10$.
III The Mpemba effect
@@
To interpret the Mpemba effect, let us consider two systems: first one $A$, initially in a steady state corresponding to a hot temperature $T_{h}$ and second one $B$, initially in a different steady state corresponding to cold temperature $T_{c}$ where $T_{h}>T_{c}$. These initial steady states are denoted by $\boldsymbol{\pi}(T_{h})$ and $\boldsymbol{\pi}(T_{c})$ respectively. Now imagine that both $A$ and $B$ are quenched at once to a common bath temperature, $T_{b}$, where $T_{h}>T_{c}>T_{b}$. Eventually, both of them will relax to the common distribution $\boldsymbol{\pi}(T_{b})$ given long enough time. The Mpemba effect is said to exist if $A$ reaches the final steady state faster than $B$ during the transient/relaxation process. This relaxation can be quantified in terms of the distance function $D[\boldsymbol{\mathcal{P}}(t),\boldsymbol{\Pi}(T_{b})]$ which measures the instantaneous distance of the state-wise occupation probabilities $\boldsymbol{\mathcal{P}}(t)=(P_{1},P_{2},P_{3})^{\mathbb{T}}$ from their respective final steady state distributions $\boldsymbol{\Pi}(T_{b})=(\pi_{1},\pi_{2},\pi_{3})^{\mathbb{T}}$. There are many well-adapted measures that exist in the literature namely the norm $L_{1}(t)=\sum_{i=1}^{3}|P_{i}(t)-\pi_{i}(T_{b})|$ and Kullback-Leibler (KL) divergence $KL(t)=\sum_{i=1}^{3}P_{i}(t)\ln(P_{i}(t)/\pi_{i})$. Thus, if one has $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]>D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ initially for $T_{h}>T_{c}$ followed by $D[\boldsymbol{\Pi}(T_{h}),\boldsymbol{\Pi}(T_{b})]<D[\boldsymbol{\Pi}(T_{c}),\boldsymbol{\Pi}(T_{b})]$ at a later time, we infer that the Mpemba effect exists.
IV Activity harnessed Mpemba effect
To highlight the effect of activity, we first set $\epsilon=\gamma=0$ and consider the passive discrete three-state Markov chain [15]. The identification of the Mpemba effect in this case can often be made from the behaviour of the coefficient $|a_{2}(T,T_{b})|$ [as in Eq. (5)], namely by checking whether $|a_{2}(T_{c},T_{b})|>|a_{2}(T_{h},T_{b})|$ or by measuring the distance functions [15]. In this set-up, let us now choose a set of parameters $R_{ij}$-s such that the Mpemba effect is absent. However, as the activity is introduced by setting $\epsilon,\gamma>0$, we notice a spontaneous emergence of the Mpemba effect [see Fig. 3]. We also observe a converse effect where the introduction of activity suppresses an already existing Mpemba effect (Fig. 3). Quite distinctly, this behavior remains robust under other distance measures. To see this, we measure the relaxation in terms of two popular distance measures: $L_{1}$-norm and KL divergence. Figure 4 shows, for a particular choice of the model parameters, that both the measures infer the existence of the Mpemba effect. For convenience, we use $L_{1}$-norm in all the instances unless otherwise stated. It is worthwhile to note that the measure $|a_{2}(T,T_{b})|$ is in general not an appropriate measure to describe the Mpemba effect in systems with broken detailed balance and complex eigenspectrum. This is because such systems do not necessarily show monotonic relaxations and consequently, multiple crossings between the relaxation trajectories prepared from different initial conditions can emerge, as will be discussed later (also see [30, 35, 41]). This is in contrast with monotonic relaxations (as in the case of $\epsilon=\gamma=0$) where the contribution of the second eigenmode reflected by $|a_{2}(T,T_{b})|$ becomes crucial (see the discussion in Appendix D).